OpenStudy (anonymous):

Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose.

OpenStudy (lalaly):

x + 2 = y Using the property of equality, subtract 2 from both sides of the equal operator. x + 2 - 2 = y - 2 Now, simplify by combining like terms. x = y - 2

OpenStudy (anonymous):

thank you, you're a lifesaver (:

OpenStudy (lalaly):

ur welcome:D